# PowerInfer: Fast Large Language Model Serving with a Consumer-grade GPU

### v1

@@ 3.1 Insight-1: Power-law Activation
Figure 5: Cumulative distribution function (CDF) of neuron activation in OPT-30B and LLaMA(ReGLU)-70B. (a) CDF in a single MLP layer. (b) CDF across the entire model. The X-axis shows neuron proportion. The Y-axis represents the CDF of neuron activation.
@@
LLM inference exhibits a high degree of locality, indicating that a consistent group of neurons is frequently activated. Notwithstanding the input dependence of LLM activation sparsity, a power-law distribution is evident among activated neurons. Figure 5a reveals that in the MLP layers of OPT-30B and LLaMA (ReGLU)-70B, 26% and 43% of neurons respectively are responsible for 80% of total activations. These are termed hot-activated neurons. Conversely, the activation of the remaining 74% and 57% of neurons is input-dependent, classifying them as cold-activated neurons.
@@
This high locality is not confined to a single layer but extends throughout the model. As illustrated in Figure 5b, approximately 17% of neurons in OPT-30B and 26% in LLaMA (ReGLU)-70B are responsible for 80% of the total activations across all layers.
3.2 Insight-2: Fast In-CPU Computation
(a) MLP layer
(b) Attention layer
@@ 3.2 Insight-2: Fast In-CPU Computation
Figure 6: Comparison of execution time for load-then-execute versus direct-execute methods when 10% and 60% neuron weights of one MLP and attention layer in OPT-30B are CPU-resident. The X-axis shows input batch sizes, and the Y-axis measures execution time (ms). Load-then-execute involves transferring these neuron weights to GPU memory for computation, whereas direct-execute computes them directly on the CPU.
@@
We compared the time to load and compute 10% [^3] While Insight-1 indicates that 43% of neurons account for 80% of the total activations in a single MLP layer, it is typically found that only about 10% of its neurons are activated during an individual inference iteration. of the MLP layer and 60% of attention layer’s CPU-side neurons on the GPU versus direct CPU execution in OPT-30B. Results in Figure 6 indicate that for batch sizes under 32, the time taken to transfer the weights of these neurons and compute them on the GPU (NVIDIA RTX 4090) exceeds the time required for calculation directly on the CPU using the AVX2 vector extension.
4 PowerInfer Overview
[…]
@@ 4 PowerInfer Overview
4.1 Architecture and Workflow
Figure 7: The architecture overview and inference workflow of PowerInfer.
@@ 4.1 Architecture and Workflow
LLM Profiler and Policy Solver (Offline): This component includes an LLM profiler that collects activation data from inference processes using requests derived from general datasets (e.g., C4 [32]). It monitors neuron activation across all layers (Step ①), followed by a policy solver categorizing neurons as hot or cold. The solver aims to allocate frequently activated neurons to the GPU and others to the CPU. It uses a neuron impact metric and hardware specifications to balance the workload, using integer linear programming to maximize the GPU’s impact metric for neurons (Step ②).
@@
Figure 8: An illustrative example shows how PowerInfer calculates different neurons for one LLM layer.
4.2 Single Layer Example
@@
Both CPU and GPU then process predicted active neurons, ignoring inactive ones. The GPU computes neurons 3 and 5, while the CPU handles neuron 4. Once neuron 4’s computation is complete, its output is sent to the GPU for result integration.
5 Neuron-aware Inference Engine
[…]
5.1 Adaptive Sparsity Predictors
The online inference engine in PowerInfer reduces computational loads by only processing those neurons that are predicted to be activated. This method was also used in DejaVu [21], which advocates for training a set of fixed-size MLP predictors. Within each Transformer layer, DejaVu utilizes two separate predictors to forecast the activation of neurons in the self-attention and MLP blocks. Consequently, the inference computation is confined to neurons anticipated to be active.
However, designing effective predictors for local deployments with limited resources is challenging, balancing prediction accuracy and model size. These predictors, frequently invoked for neuron activation prediction, should be stored in GPU memory for fast access. Yet, the considerable memory requirements of numerous fixed-size predictors can encroach upon the space needed for storing LLM parameters. For example, predictors for the OPT-175B model require around 27GB of GPU memory, surpassing an NVIDIA RTX 4090 GPU’s capacity. On the other hand, naively reducing predictor size may impair accuracy; a decrease from 480MB to 320MB in predictor size dropped its accuracy from 92% to 84%, further adversely affecting the overall LLM accuracy (e.g., winogrande [35] task accuracy from 72.77% to 67.96%).
Figure 9: Correlation between predictor parameter size and layer sparsity at a guaranteed 95% accuracy level for OPT-175B. The X-axis represents sparsity, and the Y-axis represents the predictor parameter size. The bar indicates the average parameter size for the model in the corresponding sparsity, while the error bar reflects fluctuations in the predictor parameter size due to skewness within the layer.
@@ 5.1 Adaptive Sparsity Predictors
To optimize for these factors, PowerInfer designs an iterative training method for non-fixed-size predictors for each Transformer layer. The process begins by establishing a baseline model size based on the layer’s sparsity profile (Figure 9). Subsequently, the model size is iteratively adjusted, taking into account the internal activation skewness to maintain accuracy. An MLP predictor typically comprises input, hidden, and output layers. Since the dimensions of the input and output layers are determined by the Transformer layer’s structure, modifications primarily target the hidden layer. During the iterative adjustments, the hidden layer’s dimension is modified according to the observed skewness. For layers exhibiting significant skewness, the hidden layer size is reduced progressively until accuracy falls below 95%. Conversely, for layers with minimal skewness, the dimension is increased to improve accuracy. Through this approach, PowerInfer effectively limits predictor parameters to a mere 10% of the total LLM parameters.
5.2 Neuron Placement and Management
@@
Ensuring the accurate computation of these segmented neurons in their proper sequence is vital for precise results. To this end, PowerInfer creates two neuron tables, one located in the CPU and the other in the GPU memory. These tables correlate each neuron to its original position in the matrix. During the process of multiplying with an input tensor, each neuron interacts with its corresponding tensor value, guided by the mappings in the neuron tables. The additional memory required for these neuron tables is relatively insignificant, totaling only about 9MB for an LLM like OPT-175B, which needs 350GB of storage.
5.3 GPU-CPU Hybrid Execution
Given that PowerInfer processes only a limited number of neurons predicted to be active, such as less than 10% in an MLP layer, a potential method for GPU and CPU collaboration involves transferring cold-activated neuron weights from the CPU to the GPU for computation. However, as per Insight-2, the time spent transferring activated neurons to the GPU surpasses the time needed for direct computation on the CPU. Therefore, PowerInfer implements a GPU-CPU hybrid execution model, wherein both units independently compute their respective activated neurons and then combine the results on the GPU. This method effectively balances the computational workload, leveraging the strengths of each unit while reducing transfer time inefficiencies.
@@ 5.3 GPU-CPU Hybrid Execution
In scenarios where activated neurons are split between GPU and CPU, synchronization between these processing units also becomes crucial. After one unit finishes its neuron calculations, it waits for the other to merge results. As GPU neurons are activated more frequently, PowerInfer assigns merging operations to the GPU. To optimize synchronization overhead, a selective synchronization strategy is used, bypassing result synchronization when the CPU executor has no activated neurons, allowing it to proceed to subsequent blocks, thereby enhancing overall efficiency.
5.4 Neuron-aware Operator
Considering the activation sparsity in LLMs, matrix multiplication operations can bypass inactive neurons and their weights, necessitating the use of sparse operators. However, current sparse matrix multiplication tools, including state-of-the-art sparse-aware compilers like SparTA [52] and FlashLLM [45], as well as libraries like cuSPARSE [30] and Spunik [33], fall short in this regard. They either support only static compilation of sparse-aware kernels or require dynamic conversion of sparse matrices into dense formats, leading to significant performance overhead, especially with the dynamic sparsity in our scenario. Additionally, the dynamic JIT compiler PIT [51], though efficient for general sparse matrix multiplication on GPUs, is not suited for CPU-GPU hybrid execution where CPU computational capabilities are limited.
@@ 5.4 Neuron-aware Operator
Neuron-aware Operators for GPU: Despite vector-vector calculations being less efficient than matrix-vector calculations on GPU, neuron-aware operators based on vector-vector computation are advantageous when the batch size is small. They avoid unnecessary computations and memory operations associated with inactive neurons and do not need costly matrix conversions. Furthermore, these operators allow all thread blocks to concurrently check neuron activations and compute corresponding vectors if activated.
Neuron-aware Operators for CPU: Neuron-aware operators are particularly beneficial for CPUs, which generally have lower parallelism and matrix computation efficiency. The CPU executor assigns a neuron-aware operator to multiple cores, dividing neurons into smaller batches for concurrent activation checking. Each core processes only the activated neurons in its batch, optimizing vector-vector calculations with hardware vector extensions like AVX2, widely supported in modern CPUs.
6 Neuron Placement Policy
To fully unleash the computational capability of the GPU and CPU, PowerInfer’s offline component provides a placement policy to guide the allocation of each neuron to either the GPU or CPU. This policy, output by a solver, controls neuron placement within each layer, thereby defining the runtime computational workload for the respective processing units. The solver considers a range of factors, including each neuron’s activation frequency, communication overhead, and the computational capacities of the processing units, such as their memory sizes and bandwidths.
The solver defines an impact metric for each neuron to model its activation information. By integrating the neuron impacts with the capabilities of different computing units, the solver constructs an integer linear programming model to generate the optimal neuron placement.
6.1 Offline Profiling
Before determining the placement of each neuron, the offline profiler of PowerInfer needs to gather runtime inference data for each neuron. To achieve this, it deploys the LLM to handle requests generated from multiple general datasets, such as C4 [32] and Wikipedia [10]. To accurately measure activation information, the profiler inserts a monitoring kernel after each block within a Transformer layer. Additionally, it builds a neuron information table on the GPU, designed to track the activation count of each neuron.
This kernel checks whether each neuron in the layer gets activated during the inference process and, if so, increments the corresponding count in the neuron table. Once all requests have been processed, the profiler retrieves the activation data from this table and passes it to the solver.
6.2 Neuron Impact Metric
@@
$v_{i}=f_{i}\hskip 28.45274pt\forall i\in\mathbb{N}$ (1)
@@
$\mathbb{L}$ Par
All layers
$\mathbb{N}$ Par
All neurons
$\mathbb{U}$ Par
CPU and GPU
$f_{i}$ Par
Activation frequency of neuron j
$N_{i}$ Par
Neuron in layer i
$v_{i}$ Par
Neuron impact for neuron j
$M_{i}$ Par
The memory size for neuron i
$MCap_{j}$ Par
The memory size for processing unit j
$Bandwidth_{j}$ Par
The memory bandwidth for processing unit j
$T_{sync}$ Par
The time required for one synchronization between the CPU and GPU
$K$ Par
A large positive number
$a_{in}$ Var
Whether neuron n is placed on processing unit U
$T_{l}^{j}$ Var
The time for computing one neuron in layer $l$ on processing j
$C_{l}$ Var
@@
The minimum number of neurons required to be allocated on the GPU when the solver opts to split neurons in layer $l$
$y_{l}$ Var
@@
Binary auxliary variable for layer $l$ to facilitate the modeling of conditional constraints
Table 1: Terminology for ILP formulation. The Par represents the parameters gathered from the profiler or the expressions used to define constraints, none of which need to be solved by the solver. The Var refers to the constraint and objective variables that emerge from the modeling process, which need to be solved by the solver.
6.3 Modeling of Neuron Placement

### v2

@@ 3.1. Insight-1: Power-law Activation
Figure 4. Cumulative distribution function (CDF) of neuron activation in OPT-30B, LLaMA2(ReGLU)-70B and LLaMA2(SwiGLU)-70B. (a) CDF in a single MLP block. (b) CDF across the entire model. The X-axis shows neuron proportion. The Y-axis represents the CDF of neuron activation.
@@
LLM inference exhibits a high degree of locality, indicating that a consistent group of neurons is frequently activated. Notwithstanding the input dependence of LLM activation sparsity, a power-law distribution is evident among activated neurons. We profile from wikipedia (Foundation, 2024) dataset to collect the statistics of the activation sparsity about 1M tokens. Figure 4a reveals that in the MLP blocks of OPT-30B (using ReLU), LLaMA2 (ReGLU)-70B and LLaMA2 (SwiGLU)-70B, 26%, 43% and 69% of neurons respectively are responsible for 80% of total activations. This indicates that these neurons are frequently activated, which we termed as hot-activated neurons. Conversely, the activation of the remaining 74%, 57% and 31% of neurons is input-dependent, classifying them as cold-activated neurons. This distribution is particularly pronounced in ReLU-based models compared to those using other activation functions due to its higher sparsity.
@@
Furthermore, this high locality is not confined to a single MLP block but extends throughout the entire model. As illustrated in Figure 4b, approximately 17% of neurons in OPT-30B, 26% in LLaMA2 (ReGLU)-70B, and 75% in LLaMA2 (SwiGLU)-70B are responsible for 80% of the total activations across all layers. Figure 5a clearly demonstrates that in the OPT-30B model, each transformer layer contains a small portion of hot neurons, while the majority are cold neurons. Notably, in the initial 24 layers, OPT-30B exhibits exceptionally low activation, with less than 1% of neurons becoming active, resulting in a minimal presence of hot neurons within these layers.
Figure 5. Activation frequency in OPT-30B. (a) The activation frequency of neurons in different layers. The Y-axis represents the layer id. (b) The activation frequency of neurons in different tasks for 30th layer. The Y-axis represents the tasks. The X-axis shows neuron proportion.
Interestingly, the same group of hot neurons consistently remains active across various downstream tasks. To demonstrate this, we profiled neuron activation across diverse datasets, including knowledge-based, truthfulness assessment, and reasoning tasks. Figure 5b illustrates that these hot neurons are consistently activated consistently across different tasks, indicating a stable pattern of power-law activation. Our analysis revealed that there is over 90% overlap in the top 20% most frequently activated neurons across these varied domains. This remarkable consistency suggests that the power-law distribution of neuron activations is an intrinsic property of the model architecture rather than being dataset-specific.
3.2. Insight-2: Fast In-CPU Computation
@@
Figure 6. Comparison of execution time for load-then-execute versus direct-execute methods when 10% and 60% neuron weights of one MLP and attention block in OPT-30B are CPU-resident. The X-axis shows input batch sizes, and the Y-axis measures execution time (ms). Load-then-execute involves transferring these neuron weights to GPU memory for computation, whereas direct-execute computes them directly on the CPU.
@@
We compared the time to load and compute 10% [^3] While Insight-1 indicates that 43% of neurons account for 80% of the total activations in a single MLP block, it is typically found that only about 10% of its neurons are activated during an individual inference iteration. sparisty of the MLP block and 60% of attention block’s CPU-side neurons on the GPU versus direct CPU execution in OPT-30B. Results in Figure 6 indicate that for batch sizes under 32, the time taken to transfer the weights of these neurons and compute them on the GPU (NVIDIA RTX 4090) exceeds the time required for calculation directly on the CPU using the AVX2 vector extension.
4. PowerInfer Overview
We present PowerInfer, a low-latency LLM inference system deployed in a PC equipped with a single consumer-grade GPU. PowerInfer proposes a neuron-aware offloading strategy and an inference engine by fully leveraging the high locality insights described in §$\S$3. It utilizes both GPU and CPU for weight storage, accommodating LLMs of various sizes. This offloading approach, based on Insight-1, effectively exploits the power-law distribution of LLM inference. Specifically, PowerInfer preloads the GPU with weights for neurons that activate frequently, while less active neurons’ weights are kept on the CPU.
@@ 4. PowerInfer Overview
The effectiveness of PowerInfer is directly correlated with the model’s activation sparsity. ReLU-family LLMs, exhibiting over 90% sparse activations in their FFNs, are ideal candidates for PowerInfer. While LLMs with other activation functions typically show around 50% sparsity, resulting in less pronounced acceleration, PowerInfer still offers some performance gains. Encouragingly, there is a growing trend on enhancing LLM sparsity without compromising performance (Song et al., 2024c, a), or directly training LLMs with ReLU-family activations, as seen in NVIDIA Nemotron (Adler et al., 2024) and MiniTron (Sreenivas et al., 2024). This trend is expected to extend PowerInfer’s applicability across a more diverse spectrum of LLMs.
4.1. Architecture and Workflow
Figure 7. The architecture and inference workflow of PowerInfer.
@@ 4.1. Architecture and Workflow
LLM Profiler and Policy Solver (Offline): Based on Insight-1, the neuron’s activation pattern can be collected enough from general datasets, so we take an offline profiler that collects activation data from inference processes using requests derived from general datasets (e.g., C4 (Raffel et al., 2020)). It monitors neuron activation across all layers (Step ①), followed by a policy solver categorizing neurons as hot or cold. The solver aims to allocate frequently activated neurons to the GPU and others to the CPU. It uses a neuron impact metric and hardware specifications to balance the workload, using integer linear programming to maximize the GPU’s impact metric for neurons (Step ②).
@@
Figure 8. An illustrative example shows how PowerInfer calculates different neurons for one LLM layer.
4.2. Single Layer Example
@@
Both the GPU and CPU concurrently process their respective predicted active neurons, while efficiently bypassing inactive ones. Specifically, the GPU computes neurons 3 and 5, leveraging its parallel processing capabilities, while the CPU simultaneously handles neuron 4. Upon completion of neuron 4’s computation on the CPU, its output is swiftly transferred to the GPU. The GPU then performs a final integration step, combining the results from all activated neurons (3, 4, and 5) using an optimized add operator.
5. Neuron-aware Inference Engine
5.1. Adaptive Sparsity Predictors
The online inference engine in PowerInfer reduces computational loads by only processing those neurons that are predicted to be activated. This method was also used in DejaVu (Liu et al., 2023b), which advocates for training a set of fixed-size MLP predictors. Within each Transformer layer, DejaVu utilizes two separate predictors to forecast the activation of neurons in the self-attention and MLP blocks. Consequently, the inference computation is confined to neurons predicted to be active.
However, designing effective predictors for local deployments with limited resources is challenging, balancing prediction accuracy and model size. These predictors, frequently invoked for neuron activation prediction, should be stored in GPU memory for fast access. Yet, the considerable memory requirements of numerous fixed-size predictors can encroach upon the space needed for storing LLM parameters. For example, predictors for the OPT-175B model require around 27GB of GPU memory, surpassing an NVIDIA RTX 4090 GPU’s capacity. On the other hand, naively reducing predictor size impairs accuracy; a decrease from 480MB to 320MB in predictor size dropped its accuracy from 92% to 84%, further adversely affecting the overall LLM accuracy (e.g., winogrande (Sakaguchi et al., 2021) task accuracy from 72.77% to 67.96%).
Figure 9. Correlation between predictor parameter size and layer sparsity at a guaranteed 95% accuracy level for OPT-175B. The X-axis represents sparsity, and the Y-axis represents the predictor parameter size. The bar indicates the average parameter size for the model in the corresponding sparsity, while the error bar reflects fluctuations in the predictor parameter size due to skewness within the layer.
@@ 5.1. Adaptive Sparsity Predictors
To optimize for these factors, We design an iterative training method for non-fixed-size predictors for each Transformer layer. The process begins by establishing a baseline model size based on the layer’s sparsity profile (Figure 9). Subsequently, the model size is iteratively adjusted, considering the internal activation skewness to maintain accuracy. An MLP predictor typically comprises a input, a single hidden, and an output layers. Since the dimensions of the input and output layers are determined by the Transformer layer’s structure, modifications primarily target the hidden layer.
In training predictors, model perplexity on the WikiText-2 dataset served as the benchmark for the predictive accuracy. Initially, the predictor’s dimensionality was set in accordance with the sparsity level of the layer. We then engaged in an iterative process to tune the size of the hidden layer. This process was continued until the perplexity of the model with the integrated predictor closely approximated that of the baseline model, achieving a discrepancy of less than 0.1%. Through this approach, PowerInfer effectively limits predictor parameters to a mere 6% of the total LLM parameters.
5.2. Neuron Management
@@
Ensuring the accurate computation of these segmented neurons in their proper sequence is vital for precise results. To this end, PowerInfer creates two neuron tables, one located in the CPU and the other in the GPU memory. These tables correlate each neuron to its original position in the matrix. During the process of multiplying with an input tensor, each neuron interacts with its corresponding tensor value, guided by the mappings in the neuron tables. The additional memory required for these neuron tables is relatively minimal, totaling only about 9MB for an LLM like OPT-175B, which needs 350GB of storage.
5.3. GPU-CPU Hybrid Execution
PowerInfer implements a GPU-CPU hybrid execution model where both units independently compute their respective activated neurons. This approach balances the computational workload between GPU and CPU, leveraging the strengths of each unit while minimizing transfer time inefficiencies. The GPU processes preloaded hot neurons, while the CPU handles cold neurons, with results combined on the GPU.
@@ 5.3. GPU-CPU Hybrid Execution
In scenarios where activated neurons are split between GPU and CPU, synchronization between these processing units also becomes crucial. After one unit finishes its neuron calculations, it waits for the other to merge results. As GPU neurons are activated more frequently, PowerInfer assigns merging operations to the GPU.
5.4. Neuron-aware Operator
Considering the activation sparsity in LLMs, matrix multiplication operations can bypass inactive neurons and their weights, necessitating the use of sparse operators. However, current sparse matrix multiplication tools, including state-of-the-art sparse-aware compilers like SparTA (Zheng et al., 2022) and FlashLLM (Xia et al., 2023), as well as libraries like cuSPARSE (NVIDIA, 2023), fall short in this regard. They either support only static compilation of sparse-aware kernels or require dynamic conversion of sparse matrices into dense formats, leading to significant performance overhead, especially with the dynamic sparsity in our scenario. Additionally, the dynamic JIT compiler PIT (Zheng et al., 2023b), though efficient for general sparse matrix multiplication on GPUs, is not suited for CPU-GPU hybrid execution where CPU computational capabilities are limited.
@@ 5.4. Neuron-aware Operator
Neuron-aware Operators for GPU: Despite vector-vector calculations being less efficient than matrix-vector calculations on GPU, neuron-aware operators based on vector-vector computation are advantageous when the batch size is small. They avoid unnecessary computations and memory operations associated with inactive neurons and do not need costly matrix conversions. Furthermore, these operators allow all thread blocks to concurrently check neuron activations and compute corresponding vectors if activated. Notably, PowerInfer assigns independent computational tasks to different thread blocks on the GPU. Each thread block processes a distinct set of neurons, eliminating the need for synchronization or divergence across blocks.
Neuron-aware Operators for CPU: Neuron-aware operators are particularly beneficial for CPUs, which generally have lower parallelism. The CPU executor assigns a neuron-aware operator to multiple cores, dividing neurons into smaller batches for concurrent activation checking. Each core processes only the activated neurons in its batch, optimizing vector-vector calculations with hardware vector extensions like AVX2.
6. Neuron Placement Policy
To fully harness the computational power of GPUs and CPUs, a locality-aware policy is crucial for optimizing performance. If we randomly place neurons on GPU and CPU, the GPU cannot fully utilize its powerful computation capability because some hot neurons are placed on CPU. Further, simply assigning the hotest neurons to the GPU may lead to excessive data transfers between the CPU and GPU, undermining efficiency. To tackle this challenge, PowerInfer’s offline component devises a placement policy using a solver that determines the allocation of each neuron to GPU or CPU.
6.1. Offline Profiling
Before determining the placement of each neuron, it is important to profile the activation information of each neuron. Based on Insight-1 that the hot neurons in general corpus are also activated frequently across different scenarios, we achieve the profile with an offline profiler, which deploys the LLM to handle requests generated from multiple general datasets, such as C4 (Raffel et al., 2020) and Wikipedia (Foundation, 2024). To accurately measure activation information, the profiler inserts a monitoring kernel after each block within a Transformer layer. Additionally, it builds a neuron information table on the GPU, designed to track the activation count of each neuron.
6.2. Neuron Impact Metric
@@
(1) $v_{i}=f_{i}\hskip 28.45274pt\forall i\in\mathbb{N}$
Table 2. Terminology for ILP formulation. The Par represents the parameters gathered from the profiler or the expressions used to define constraints, none of which need to be solved by the solver. The Var refers to the constraint and objective variables that emerge from the modeling process, which need to be solved by the solver.
@@
$\mathbb{L}$ Par All layers
$\mathbb{N}$ Par All neurons
$\mathbb{U}$ Par CPU and GPU
$f_{i}$ Par Activation frequency of neuron i
$N_{i}$ Par Neuron in layer i
$v_{i}$ Par Neuron impact for neuron i
$M_{i}$ Par The memory size for neuron i
$MCap_{j}$ Par The memory size for processing unit j
$Bandwidth_{j}$ Par The memory bandwidth for processing unit j
$T_{sync}$ Par The time required for one synchronization between the CPU and GPU
$K$ Par A large positive number
$a_{in}$ Var Whether neuron n is placed on processing unit i
$T_{l}^{j}$ Var The time for computing one neuron in layer $l$ on processing unit j
@@
$C_{l}$ Var The minimum number of neurons required to be allocated on the GPU when the solver opts to split neurons in layer $l$
@@
$y_{l}$ Var Binary auxliary variable for layer $l$ to facilitate the modeling of conditional constraints
6.3. Modeling of Neuron Placement